Quise14:

x = 1/3y written in general form

Sherixn:

General form is getting y by itself

Sherixn:

So try and divide both sides by 1/3

Joe348:

Multiply both sides by 3 \[3x-y=0\]

Sherixn:

@joe348 wrote:
Multiply both sides by 3 \[3x-y=0\]
Yes this also works (:

jhonyy9:

\[x = \frac{ 1 }{ 3 }y\] do you think in this way ? in this case multiplie both sides by 3 like a first step